This question has multiple parts. Work all the parts to get the most points. A titration of 21.7 mL of a solution of the weak just b please thank you.
Answer #1

We have 21.7 mL of aniline to which 24.35 mL(=0.02435 L) of 0.175 M HCl was added to reach the equivalence point.

The reaction that happens is

C6H5NH2(aq) + H2O (from HCI) + CHEN H(aq) + H200

Now, number of moles of HCl added is

0.175 mol x 0.02435 L 0.00426 mol 1 L

Since this addition resulted in reaching the equivalence point, the starting moles of aniline must be 0.00426 mol.

Hence, in total we will have 0.00426 mol of C6H5NH, after reaction.

Total volume of the solution = 21.7 mL + 24.35 mL = 46.05 mL = 0.04605 L

Hence, the concentration of C6H5NH, formed is

0.00426 mol C6H5NH] = - 0.04605 L 0.0925 M

Note that C6H5NH, is an weak acid that will dissociate as follows:

C6H5NH3(aq) + H200 + C6H5NH2(aq) + H30

Given that Kb of aniline is K; = 4.0 x 10-10

We can calculate the Ka of C6H5NH, as follows

KaxKb = 1.0 x 10-14 1.0 x 10-14 ha= 4.0 X 10-1 4.0 x 10-10 = 2.5 x 10-5

Now, we can create the following ICE chart to calculate the equilibrium number of moles of each species present according to the above reaction.

C6H5NH] C6H5NH2 40*НІ
Initial, M 0.0925 0 0
Change, M -x +x +x
Equilibrium, M 0.0925-x x x

Hence, we can write the Ka expression for C6H5NH, dissociation from the above ICE chart as follows:

Ka= H30+][C6H5NH2] XXX CH5NH,110,0925 --- =2.5 x 10-5 = 2 = 2.5 x 10-5 (0.0925 – ) 1.51 x 10-3 M.

Hence, the equilibrium concentrations are

H30+1 = 1 = 1.51 x 10-3 M = 0.00151 M

C6H5NH] = 0.0925 M - r = 0.0925 M – 1.51 x 10-3 M C6H5NH10.0910 M = 9.10 x 10-2 M

Now, from H3O+ concentration, we can find the OH- concentration as follows:

[OH-][H30+1 = 1.00 x 10-14 1.00 x 10-14 → [OH-]= 1.51 x 10-3 = 6.63 x 10-12 M

Hence, the equilibrium concentrations at equivalence point are

[H3O+] = 1.51 x 10-3 M

OH-] = 6.63 x 10-12 M

[C6H5NH2) = 9.10 x 10-2 M